OpenStudy (anonymous):

solve for a, b, c a+b=5 (c^2)+(a^2)=(3^2) (b^2)+(c^2)=(4^2) (3^2)+(4^2)=(a+b)

OpenStudy (kkutie7):

lets rewrite this so it looks a little cleaner \[a+b=5\] \[c^{2}+a^{2}=3^{2}\] \[b^{2}+c^{2}=4^{2}\] \[3^{2}+4^{2}=a+b\]

OpenStudy (kkutie7):

ok so we have a lot to work with here. first thing you should look for is something that you can re-write to give you the answer to either a, b, or c

OpenStudy (kkutie7):

I'm going to mess with the very last equation ok this should just equal five but lets check and make sure \[3^{2}+4^{2}=a+b\] \[9+16=25\] this doesn't make sense =/ maybe we should just ignore this and look at the others?

OpenStudy (anonymous):

 This is what the problem actually looked like, if that helps at all